# Liquidity fragmentation on decentralized exchanges

### v5

@@ 4.1 Liquidity supply on high- and low-fee pools
Consistent with Figure 7, most of the capital deployed to provide liquidity for a given pair is locked in high-fee pools. At the same time, low-fee pools attract much larger trading volume. Models (1) and (5) show that the average low-fee pool attracts 39.5% of liquidity supply for the average pair (that is, equal to $\nicefrac{{(100-20.92)}}{{2}}$) while it executes 62% (i.e., $\nicefrac{{(100+24.62)}}{{2}}$) of the total trading volume. At a first glance, it would seem that a majority of capital on decentralized exchanges is inefficiently deployed in pools with low execution probability. We will show that, in line with our model, the difference is driven by heterogeneous liquidity cycles across pools, leading to the formation of LP clienteles.
@@
On the liquidity demand side, trades on the low-fee pool are 25.91% smaller (Model 2), consistent with Prediction 3. However, the low-fee pool executes almost three times the number of trades (i.e., trade count is 177% higher from Model 4) and has 143% higher volume than the high-fee pool (Model 3). On average, liquidity providers on low-fee pools earn 2.03 basis points higher revenue than their counterparts on high-fee pools (Model 6), indicating significant positive returns resulting from economies of scale.
@@
How does variation in fixed transaction costs impact the gap between individual order size across pools? We find that increasing the gas price by one standard deviation leads to higher liquidity deposits on both the low- and the high-fee pools (14.2% and 30.1% higher, respectively). [^9] The relative effects are computed as $\nicefrac{{0.37}}{{(1.88+0.73)}}=13.8\%$ for low pools and $\nicefrac{{0.58}}{{1.88}}=30.85\%$ for high-fee pools, respectively. The result supports Prediction 4 of the model. Our theoretical framework implies that a larger gas price leads to some (marginal) LPs switching from the low- to the high-fee pool. The switching LPs have low capital endowments relative to their low-fee pool peers, but higher than LPs on the high-fee pool. Therefore, the gas-driven reshuffle of liquidity leads to a higher average endowment on both high- and low-fee pools. Consistent with the model, a higher gas price leads to fewer active liquidity providers, particularly on low-fee pools. Specifically, a one-standard increase in gas costs leads to a significant decrease in the number of LP wallets interacting daily with low- and high-fee pools, respectively (Model 5).
While a higher gas price is correlated with a shift in liquidity supply, it has a muted impact on liquidity demand on low-fee pools. A higher gas cost is associated with 6% larger trades (Model 2), likely as traders aim to achieve better economies of scale. At the same time, the number of trades on the low-fee pool drops by 5.1% (Model 4) – since small traders might be driven out of the market. The net of gas prices effect on aggregate volume on the low-fee pool is small and not statistically significant (Model 3). The result matches our model assumption that the aggregate order flow on low-fee pool is not sensitive to gas prices. [^10] Formally, one could extend the model to assume that small LTs arrive at the market at rate $\tilde{\theta}\left(\Gamma\right)\mathop{}\!\mathrm{d}t$ and demand $f\left(\Gamma\right)$ units each, where $\tilde{\theta}\left(\Gamma\right)$ decreases in $\Gamma$ and $f\left(\Gamma\right)$ increases in $\Gamma$ such that $\tilde{\theta}\left(\Gamma\right)f\left(\Gamma\right)=\theta$.
On the high-fee pool, a higher gas price is also associated with a higher trade size, but a much smaller relative increase in the number of trades. The implication is that large-size traders who typically route orders to the high-fee pool are unlikely to be deterred by an increase in fixed costs, while they do adjust quantities to achieve better economies of scale. In the context of our model, the large LT arrival rate $\lambda$ does not depend on the gas price $\Gamma$.
@@
4.2 Liquidity cycles on high- and low-fee pools
Next, we test Predictions 5 and 6 on the duration of liquidity cycles on fragmented pools. Since the descriptive statistics in Table 1 suggest that LPs manage their positions over multiple days, we cannot accurately measure liquidity cycles in a pool-day panel. Instead, we use intraday data on liquidity events (either mints or burns) to measure the duration between two consecutive opposite-sign interactions by the same Ethereum wallet with a liquidity pool: either a mint followed by a burn, or vice-versa.
To complement our previous analysis, we additionally control for whether each liquidity position is out of range (i.e., the price range set by the LP does not straddle the current price and therefore the LP does not earn fees). We further introduce wallet fixed effects to soak up variation in reaction times across traders.
Table 5 presents the results. Liquidity updates on decentralized exchanges are very infrequent, as times elapsed between consecutive interactions are measured in days or even weeks. In line with Prediction 5, we find evidence for shorter liquidity cycles on low-fee pools. The average time between consecutive mint and burn orders is 20.06% shorter on the low-fee pool (from Model 1, the relative difference is 99.74 hours/497.18 hours).
Liquidity cycles are in part driven by fixed Blockchain transaction costs. A one standard deviation increase in gas prices speeds up the liquidity cycle on low-fee pools by a further 15.80 hours, a result which supports the intuition behind Prediction 6. When the gas price spikes, the liquidity supply on the low-fee pool decreases at a higher rate than the liquidity demand. As a result, liquidity deposits deplete faster (i.e., they move outside the fee-earning range), triggering the need for more frequent updates. At the same time, a higher gas fee speeds up the liquidity cycle on the high-fee pool as well. The result is consistent with small liquidity providers on the high-fee pool being crowded out by the high fixed costs, leading to a lower liquidity supply.
@@ 4.2 Liquidity cycles on high- and low-fee pools
Mint-burn time Burn-mint time
@@
$d_{\text{low-fee}}$99.74*** -100.17*** -104.09*** -157.95*** -159.71*** -159.69***
(-8.86) (-8.94) (-9.22) (-10.59) (-10.81) (-10.80)
Gas price $\times$ $d_{\text{low-fee}}$16.65** -15.41* -15.80** -11.29 2.95 2.94
(-2.13) (-1.98) (-2.02) (-1.65) (0.40) (0.39)
Gas price $\times$ $d_{\text{high-fee}}$14.44** -13.42* -13.98* -10.52* 1.96 1.95
(-2.04) (-1.89) (-1.96) (-1.69) (0.32) (0.32)
Volume -5.87 -7.45 -24.84*** -24.82***
(-1.15) (-1.41) (-4.10) (-4.09)
Total value locked -53.17* -51.72* -12.71 -12.71
(-1.70) (-1.66) (-0.52) (-0.52)
Volatility -2.11*** -2.26*** -2.99*** -2.98***
(-2.75) (-2.86) (-3.36) (-3.36)
Position out-of-range 37.09*** -1.53
(6.43) (-0.22)
Constant 497.18*** 497.00*** 479.22*** 248.00*** 250.13*** 250.47***
@@
The reaction time of a trader may depend on capital constraints, network congestion, and other confounding factors that can correlate with gas prices. As a robustness check, we repeat the analysis above with burn-to-mint times as the dependent variables. The burn-to-mint time measures the speed at which LPs deposit liquidity at updated prices after removing (out-of-range) positions, and should not depend on the rate at which liquidity is consumed. Consistent with the theory, we find no significant relationship between gas price and the burn-to-mint duration (Models 4 through 6).
@@ 4.3 Adverse selection costs across low- and high-fee pools
In our model, the first-order friction driving liquidity fragmentation is the fixed gas cost, which leads to heterogeneous rebalancing costs for small and large market makers respectively. Thus, pool fees in our setup simply offset liquidity providers’ costs to manage their position. An alternative channel driving fragmentation might be adverse selection: The rationale is that liquidity providers gravitate towards high-fee pools as they provide superior protection against informed traders (following the arguments in Milionis, Moallemi, and Roughgarden, 2023).
For example, consider a token pair traded in the 5 (low) and 30 (high) basis point fee pools. If an arbitrage opportunity is worth 50 basis points and the relative gas cost is 15 basis points, arbitrageurs would optimally execute swaps on both pools since the trade value surpasses total cost, but obtain lower profits on the high-fee pool. However, if gas cost rises to 25 basis points, arbitrageurs would only trade in the low fee pool. Thus, large gas costs might make the high-fee pool more appealing to liquidity providers, as it offers added protection against adverse selection. The mechanism could provide an alternative explanation for the result in Table 2 showing that market makers shift to high-fee pools when gas prices rise. In this section, we investigate differences in adverse selection across low- and high-fee pools, as well as the role of asymmetric information costs in driving liquidity fragmentation.
Our main metric for informational costs is the loss-versus-rebalancing (LVR), as defined in Milinois, Moallemi, Roughgarden, and Zhang (2023). The measure is equivalent to the adverse selection component of the bid-ask spread in equity markets. To calculate it, for each swap $j$ exchanging $\Delta x_{j}$ for $\Delta y_{j}$ in a pool with assets $x$ and $y$, we use:
@@
LVR (1h horizon) LVR (after swap) Impermanent loss TVL share Volume share CEX price deviation
@@
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
$d_{\text{low-fee}}$ 6.39*** 6.39*** 29.78*** 29.67*** 1.08*** 1.13*** -21.00*** 24.43*** 0.06 -0.00
(16.57) (17.05) (14.86) (14.95) (5.72) (6.18) (-26.45) (19.79) (1.51) (-0.06)
Gas price $\times$ $d_{\text{low-fee}}$0.75** 3.51** -0.01 -4.41*** -6.16*** 0.09
(-2.05) (2.10) (-0.05) (-6.52) (-5.64) (0.95)
Gas price $\times$ $d_{\text{low-fee}}$$\times$ LVR -0.38* -0.48 -0.10*
(-1.88) (-1.45) (-1.68)
Gas price 2.61*** 6.16** 3.71*** 1.53*** 2.82*** -0.09
(2.74) (2.53) (3.76) (4.73) (5.23) (-0.83)
Gas price $\times$ LVR 0.59*** 0.56*** 0.11**
(4.26) (2.80) (2.17)
Loss-versus-rebalancing (LVR) 0.41*** 0.67*** 0.11***
(2.72) (3.86) (6.05)
Volume 3.22*** 8.67*** 1.81*** -0.02 -0.30*** 0.18***
(8.15) (6.61) (6.22) (-0.56) (-3.18) (3.82)
Total value locked 0.53 -2.12 1.93 -0.13 0.46 -0.33***
(0.14) (-0.34) (0.74) (-0.52) (0.77) (-4.00)
Volatility 1.85*** 4.23*** 6.69** -0.36 -1.27*** 0.77***
(2.87) (3.17) (2.61) (-1.16) (-3.13) (3.16)
Constant 7.85*** 7.86*** 8.88*** 8.97*** 7.37*** 7.51*** 60.22*** 41.59*** 0.60*** 0.64***
(40.71) (36.89) (8.87) (8.88) (77.84) (61.32) (162.17) (71.02) (30.71) (38.64)
Pair FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Observations 40,302 40,288 40,302 40,288 40,250 40,248 40,288 36,059 5,207 5,207
R-squared 0.14 0.15 0.09 0.10 0.09 0.11 0.10 0.13 0.10 0.11
@@
Table 6 presents our empirical results. In line with Milionis, Moallemi, and Roughgarden (2023), all price impact measures — that is, the immediate and one-hour horizon LVR and the impermanent loss — are significantly larger in low-fee pools. This indicates that a higher liquidity fee indeed acts as barrier to arbitrageurs. Specifically, the permanent price impact, measured by the one-hour horizon LVR, is 6.85 basis points or 81% larger in low-fee pools. The total price impact, represented by the after-swap LVR metric, is 3.5 times larger in low-fee pools. The wide gap between permanent and total price impact highlights the substantially higher volume of uninformed trading in low-fee pools. Further, an increase in gas prices leads to a 3.51 wider gap in total price impact but a 0.75 bps narrower gap in permanent price impact between the high- and low-fee pools. The result suggests that a higher gas price primarily discourages uninformed traders, rather than arbitrageurs, from trading on high-fee pools.
The findings in Table 2 reveal that a higher gas cost translates to a lower liquidity and volume market share for low-fee pools. We attribute this effect primarily to the fixed costs associated with re-balancing. To examine whether the impact of gas prices on market share is driven by adverse selection, in Models (7) and (8) we add an interaction between our permanent price impact measure (one-hour horizon LVR), a dummy variable for low-fee pools, and the standardized gas price. The analysis reveals that adverse selection contributes only 8% (computed as $\nicefrac{{0.38}}{{4.41+0.38}}$) to the effect of gas prices on liquidity market shares. Moreover, this coefficient is only marginally statistically significant. A similar pattern is observed for trading volume shares in Model 8.
@@
Figure 10 graphically illustrates the result, contrasting permanent price impact measures against the liquidity yield, as calculated in equation (15). Notably, before accounting for gas costs, we observe that liquidity providers in low-fee pools experience losses on average: the average daily permanent price impact in these pools is 14.21 basis points, which exceeds the fee revenue of 11.71 bps. In contrast, liquidity providers in high-fee pools approximately break even before considering gas costs: the fee revenue amounts to 9.69 bps, which is slightly higher than the permanent price impact of 7.85 bps.

### v7

@@ 4.1 Liquidity supply on high- and low-fee pools
Consistent with Figure 7, we show in Table 3 that most of the capital deployed to provide liquidity for a given pair is locked in high-fee pools. At the same time, low-fee pools attract much larger trading volume. Models (1) and (5) show that the average low-fee pool attracts 39.5% of liquidity supply for the average pair (that is, equal to $\nicefrac{{(100-20.92)}}{{2}}$) while it executes 62% (i.e., $\nicefrac{{(100+24.62)}}{{2}}$) of the total trading volume. At a first glance, it would seem that a majority of capital on decentralized exchanges is inefficiently deployed in pools with low execution probability. We will show that, in line with our model, the difference is driven by heterogeneous rebalancing costs across pools, leading to the formation of LP clienteles.
@@
This table reports the coefficients of the following regression:
$\text{MarketShare}_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{low-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
@@
This table reports the coefficients of the following regression:
$y_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
@@
On the liquidity demand side, trades on the low-fee pool are 25.91% smaller (Model 2), consistent with Prediction 3. However, the low-fee pool executes almost three times the number of trades (i.e., trade count is 177% higher from Model 4) and has 143% higher volume than the high-fee pool (Model 3).
Next, in line with Prediction 4, low-fee pools generate a higher liquidity yield. On average, liquidity providers on low-fee pools earn 2.03 basis points higher revenue than their counterparts on high-fee pools (Model 6), indicating significant positive returns resulting from economies of scale.
@@
How does variation in fixed transaction costs impact the gap between individual order size across pools? We find that increasing the gas price by one standard deviation leads to higher liquidity deposits on both the low- and the high-fee pools (14.2% and 30.1% higher, respectively). [^9] The relative effects are computed as $\nicefrac{{0.37}}{{(1.88+0.73)}}=13.8\%$ for low pools and $\nicefrac{{0.58}}{{1.88}}=30.85\%$ for high-fee pools, respectively. The result supports Prediction 5 of the model. Our theoretical framework implies that a larger gas price leads to some (marginal) LPs switching from the low- to the high-fee pool. The switching LPs have low capital endowments relative to their low-fee pool peers, but higher than LPs on the high-fee pool. Therefore, the gas-driven reshuffle of liquidity leads to a higher average endowment on both high- and low-fee pools. Consistent with the model, a higher gas price leads to fewer active liquidity providers, particularly on low-fee pools. Specifically, a one-standard increase in gas costs leads to a significant decrease in the number of LP wallets interacting daily with low- and high-fee pools, respectively (Model 5).
While a higher gas price is correlated with a shift in liquidity supply, it has a muted impact on liquidity demand on low-fee pools. A higher gas cost is associated with 7.6% larger trades (Model 2), likely as traders aim to achieve better economies of scale. At the same time, the number of trades on the low-fee pool drops by 19.7% (Model 4) – since small traders might be driven out of the market. The net of gas prices effect on aggregate volume on the low-fee pool is small and not statistically significant (Model 3). The result matches our model assumption that the aggregate order flow on low-fee pool is not sensitive to gas prices.
On the high-fee pool, a higher gas price is also associated with a higher trade size, but also an increase in traded volume. As gas prices rise, liquidity providers switch from low- to high-fee pools. The outcome is greater depth and reduced price impact for liquidity demanders on high fee pools, which leads to higher trading volume.
@@
This table reports the coefficients of the following regression:
$y_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
@@
4.2 Re-balancing activity on high- and low-fee pools
Next, we test Prediction 6 on the duration of liquidity re-balancing cycles on fragmented pools. Since the descriptive statistics in Table 2 suggest that LPs manage their positions over multiple days, we cannot accurately measure liquidity cycles in a pool-day panel. Instead, we use intraday data on liquidity events (either mints or burns) to measure the duration between two consecutive opposite-sign interactions by the same Ethereum wallet with a liquidity pool: either a mint followed by a burn, or vice-versa.
To ensure consistency with the model described in Section 2, we conduct our analysis on the entire sample as well as on a sub-sample focused solely on re-balancing events where the liquidity position falls out of range (i.e., the price range set by the LP does not straddle the current price and therefore the LP does not earn fees). We further introduce wallet fixed effects to soak up variation in reaction times across traders, and winsorize the liquidity cycle duration at the 1% level to mitigate the influence of extreme values.
Table 6 presents the results. Liquidity updates on decentralized exchanges are very infrequent, as times elapsed between consecutive interactions are measured in days or even weeks. In line with Prediction 6, we find evidence for shorter liquidity cycles on low-fee pools. The average time between consecutive mint and burn orders is 22.05% shorter on the low-fee pool (from Model 2, the relative difference is 112.42 hours/509.19 hours).
We repeat the analysis above with burn-to-mint times as the dependent variables. The burn-to-mint time measures the speed at which LPs deposit liquidity at updated prices after removing (out-of-range) positions. Our findings reveal that LPs in low-fee pools replenish liquidity 63% faster than those in high-fee pools. This supports the notion that LPs in low-fee environments are larger, more sophisticated market participants.
@@ 4.2 Re-balancing activity on high- and low-fee pools
This table reports the coefficients of the following regression:
$y_{ijtk}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
@@
Mint-burn time (hours) Burn-mint time (hours)
Out-of-range positions Full sample
@@
$d_{\text{low-fee}}$110.94*** -112.42*** -99.74*** -100.17*** -157.95*** -159.71***
(-7.49) (-7.69) (-8.86) (-8.94) (-10.59) (-10.81)
Gas price $\times$ $d_{\text{low-fee}}$14.27 -6.54 -16.65** -15.41* -11.29 2.95
(-1.49) (-0.68) (-2.13) (-1.98) (-1.65) (0.40)
Gas price $\times$ $d_{\text{high-fee}}$19.57** -12.83 -14.44** -13.42* -10.52* 1.96
(-2.34) (-1.57) (-2.04) (-1.89) (-1.69) (0.32)
Volume -16.71*** -5.87 -24.84***
(-3.24) (-1.15) (-4.10)
Total value locked -35.14 -53.17* -12.71
(-1.05) (-1.70) (-0.52)
Volatility -3.48** -2.11*** -2.99***
(-2.49) (-2.75) (-3.36)
Constant 509.19*** 509.66*** 497.18*** 497.00*** 248.00*** 250.13***
@@ 4.3 Adverse selection costs across low- and high-fee pools
Finally, we test Prediction 7 of our model, which states that LP on the low-fee pool face higher adverse selection costs. Our main metric for informational costs is the loss-versus-rebalancing (LVR), as defined in Milinois, Moallemi, Roughgarden, and Zhang (2023). The measure is equivalent to the adverse selection component of the bid-ask spread in equity markets. To calculate it, for each swap $j$ exchanging $\Delta x_{j}$ for $\Delta y_{j}$ in a pool with assets $x$ and $y$, we use:
@@
LVR (1h horizon) LVR (after swap) Impermanent loss CEX price deviation
@@
(1) (2) (3) (4) (5) (6) (7) (8)
$d_{\text{low-fee}}$ 6.39*** 6.39*** 29.78*** 29.67*** 1.08*** 1.13*** 0.06 0.04
(16.57) (17.05) (14.86) (14.95) (5.72) (6.18) (1.51) (1.33)
Gas price $\times$ $d_{\text{low-fee}}$0.75** 3.51** -0.01 0.08
(-2.05) (2.10) (-0.05) (1.09)
Gas price 2.61*** 6.16** 3.71*** -0.03
(2.74) (2.53) (3.76) (-0.28)
Volume 3.22*** 8.67*** 1.81*** 0.22***
(8.15) (6.61) (6.22) (4.74)
Total value locked 0.53 -2.12 1.93 -0.39***
(0.14) (-0.34) (0.74) (-4.05)
Volatility 1.85*** 4.23*** 6.69** 1.04***
(2.87) (3.17) (2.61) (3.51)
Constant 7.85*** 7.86*** 8.88*** 8.97*** 7.37*** 7.51*** 0.60*** 0.67***
(40.71) (36.89) (8.87) (8.88) (77.84) (61.32) (30.71) (35.75)
Pair FE Yes Yes Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes Yes Yes
Observations 40,302 40,288 40,302 40,288 40,250 40,248 5,207 5,207
R-squared 0.14 0.15 0.09 0.10 0.09 0.11 0.10 0.11
@@
Table 7 presents our empirical results. In line with Milionis, Moallemi, and Roughgarden (2023), all price impact measures — that is, the immediate and one-hour horizon LVR and the impermanent loss — are significantly larger in low-fee pools. This indicates that a higher liquidity fee indeed acts as barrier to arbitrageurs. Specifically, the permanent price impact, measured by the one-hour horizon LVR, is 6.39 basis points or 81% larger in low-fee pools. The total price impact, represented by the after-swap LVR metric, is 3.5 times larger in low-fee pools. The wide gap between permanent and total price impact highlights the substantially higher volume of uninformed trading in low-fee pools. Our secondary measure of adverse selection, the impermanent loss at 5% around the current price, is also 15% higher on low- than high-fee pools.
We note that an increase in gas prices leads to a 3.51 wider gap in total price impact but a 0.75 bps narrower gap in permanent price impact between the high- and low-fee pools. The result suggests that a higher gas price primarily discourages uninformed traders, rather than arbitrageurs, from trading on high-fee pools.
@@
Figure 11 graphically illustrates the result, contrasting permanent price impact measures against the liquidity yield, as calculated in equation (20). Notably, before accounting for gas costs, we observe that liquidity providers in low-fee pools experience losses on average: the average daily permanent price impact in these pools is 14.21 basis points, which exceeds the fee revenue of 11.71 bps. In contrast, liquidity providers in high-fee pools approximately break even before considering gas costs: the fee revenue amounts to 9.69 bps, which is slightly higher than the permanent price impact of 7.85 bps. One should keep in mind, however, that the magnitude of losses from adverse selection depends on the horizon at which we measure the loss-versus-rebalancing.